# Long Grounded Thoughts: Distilling Compositional Visual Reasoning Chains at Scale

### v1

@@ 6 Conclusion
We introduced a scalable framework for synthesizing vision-centric reasoning data that unifies scale, complexity, and cognitive richness, producing 1M+ high-quality data. Our pipeline leverages grounded metadata and composition hardening to create diverse, verifiable tasks that go well beyond prior math-centric efforts. By coupling VLMs with reasoning LLMs, we show how frontier models can be “given eyes” to distill rich cognitive behaviors while maintaining scalability. Our experiments demonstrate that this data not only advances open VLMs on vision-centric tasks, but also transfers across modalities and out of distribution data.
7 Limitations
While our models show consistent gains across image MCQ benchmarks, several limitations remain. First, instruction-following on non-MCQ formats is weaker, suggesting that mixing in more open-ended data may be necessary for real deployment. Second, despite filtering, some synthesized MCQs remain vague or ambiguous; stronger rubric-based judging with LLMs could help, though at higher cost. Finally, our image pool is limited to DOCCI, and our approach relies on high-quality captions.
Reproducibility statement
To ensure the reproducibility of our work, we provide comprehensive details of our proposed model, including model specifications, data generation pipeline, pre-processing steps, training setups in the main paper and supplementary materials, along with the specific hyperparameter configurations and the prompts for data generation. The artifacts of this work such as datasets and models are planned to be publicly released upon publication, to facilitate future works in visual reasoning.
@@ A.1 Stage 1: Object-centric Problem Generation
Object-oriented metadata generation. Our data generation pipeline is composed of two primary stages: a scalable data generation stage focused on quantity and a compositional hardening stage focused on complexity. The first stage is illustrated in Figure 6. Our pipeline begins by leveraging a dense captioning model to produce a detailed textual description of the image. Concurrently, we process the image with Grounded SAM-2 (ren2024grounded) to extract a rich set of open-vocabulary bounding boxes and corresponding object tags, which we term Object Metadata. This metadata, including labels and coordinates (e.g., bag ($0.29,0.11,0.39,0.32$)), is integrated with the dense caption into a structured system prompt. This prompt guides a large language model (LLM Gen) to create object-centric MCQs that are inherently focused on specific visual elements. On average, each DOCCI image contains a median of 10.7 bounding tags, determined after applying a confidence cutoff of 0.9 from the detection source variable of the Florence-2 feature extractor (xiao2024florence), a 0.7B vision encoder–decoder model, within the Grounded SAM-2 bounding box generation pipeline. To stabilize the bounding box–grounded MCQ generation process in the Stage 1 pipeline over diverse object-context, we constrain the maximum number of same-category instances per image to 9 (e.g., up to nine “tree” bounding box tags in a visual scene image).
Semantic filtering for question diversity. To ensure the quality and diversity of the generated data, we employ a two-step filtering process. First, an Qwen3-30B-A3B-Instruct-2507 (yang2025qwen3) based LLM verifier assesses the factual correctness and logical soundness of each generated MCQ. Second, a semantic similarity filtering step is applied to discard redundant or overly similar questions, thereby enhancing the diversity of the final dataset.
For the semantic filtering, we represent each MCQ $i$ by a question stem embedding $q_{i}$, a selected answer embedding $a_{i}$, and an optional set of category tags $c_{i}$. Embeddings are derived from all-MiniLM-L6-v2, a 22M text embedding model in SentenceTransformer (reimers2019sentence), after standard text normalization. We define a composite similarity score between two MCQs, $i$ and $j$, as a weighted combination:
@@
A.2 Vision-Centric Prompt Templates
When the previous LLM based text-to-text MCQs generation frameworks (liao2025longperceptualthoughts) are highly relied on the image caption quality of the text, injection visual-grounded information has not been well explored. We carefully design a coordinate information guided prompt for the vision-centric MCQ generation with one main task instruction prompt, one core principle, and four auxiliary guidelines as shown in  A.2. We will open source the full data generation pipeline with the exact prompt instruction.
@@ A.2 Vision-Centric Prompt Templates
Prompt Task At the start of the process, we add a “computer vision expert” role-playing instruction; therefore, the text-LLM can better interpret bounding box coordinates and image tags while generating questions. In our experiments, using this “computer vision expert” instruction led to a noticeable improvement: the instruction-following F1 score rose from $96.3$ to $99.2$ during Stage 1 data generation. Similar role-playing (zhang2025automated), self-reflection (shinn2023reflexion), and task-activating prompting (yang2023generative) approaches have also been explored in recent work on domain-specific MCQ generation, such as in medical and speech understanding.
@@
Core Principle. To strengthen the model’s ability to reason over complex spatial and visual information in MCQs, we perform an error analysis on MTVQA (tang2024mtvqa) and MegaBench (chen2024mega) benchmarks that do not overlap with our test set, using the QwenVL-2.5 report as reference. The analysis reveals that most failure cases stem from two sources: 55.1% misunderstanding spatial relationships (e.g., identifying where an object mentioned in the question is located in the image) and 32.2% misinterpreting contextual information (e.g., inferring the function or color of the subject). Motivated by these observations, we establish four vision-centric principles to guide question construction, leveraging dense captions, bounding box labels, and coordinate information.
@@
In summary, LGT employs a carefully designed prompt instruction comprising $1,987$ tokens, which is substantially longer than the simpler vision-instruction prompt of $419$ tokens used in  liao2025longperceptualthoughts. The full prompt is provided in detail below.
@@
You must generate exactly {{ num_questions | default(4) }} questions. q Questions MUST NOT disclose bounding box coordinates or specific object labels. Use generic terms like "the object," "the item," or "the element" instead of "{{ bbox_label }}". Answers MUST INCLUDE the exact object label and integer coordinates in the specified format: "{{ bbox_label }}, [{{ bbox_coordinates[0]|round|int }}, {{ bbox_coordinates[1]|round|int }}, ...], [specific answer]". t Questions must focus on visual properties or relationships of the object within the specified bounded region, requiring careful inspection.
A.3 Stage 2: Compositional Problem Generation
@@
Simple CoT VLM distillation and Thought Expansion A core challenge in synthesizing reasoning traces is that open-weight VLMs often lack the rich, non-linear cognitive behaviors (e.g., subgoal setting, backtracking) seen in frontier LLMs, while human annotation is prohibitively expensive at scale. To address this, we adopt the two-step distillation process from LPT, as illustrated in Figure 1 (bottom). The prompt templates used for this distillation process are adapted from a baseline in liao2025longperceptualthoughts.
@@
A.4 Local Qwen CoT Verification Prompt
@@
A.5 Additional Experimental Results
A.5.1 Multimodal Audio Understanding and Omni Model
We emphasize our data is completely vision-centric. Here, we conduct experiments on out-of-domain (OOD) complex audio reasoning tasks using the Qwen2.5-Omni-7B-Instruct (xu2025qwen2) model on the Multimodal Audio Understanding (MMAU) benchmark (sakshi2024mmau). Our methodology leverages the modular architecture of Qwen-Omni, which separates the reasoning component (“thinker”) from the answer generation component (“talker”). We first fine-tune only the “thinker” dense module using our synthesized data. We keep our modality-specific audio encoder frozen when tuning the backbone dense model. After this stage, the fine-tuned “thinker” is merged back with the original, pre-trained “talker” module. This strategy aims to enhance the model’s intrinsic reasoning capabilities without directly altering the audio-specific knowledge contained within the “talker” module, thereby mitigating catastrophic forgetting.
@@ A.5.1 Multimodal Audio Understanding and Omni Model
Training on 500k LGTs already pushes the model’s performance to 71.80, outperforming the other leading omni-models. By scaling to 1M LGTs, our model achieves a new state-of-the-art average score of 72.32 among the tested configurations. These gains are driven by significant improvements in non-speech acoustic reasoning, with the MMAU-Sound score rising from 76.77 to 78.30 and the MMAU-Music score increasing from 67.33 to 70.20. This result underscores that enhancing core reasoning abilities with our vision-centric data can positively transfer to and improve performance on OOD audio tasks. This complex tracing of learning benefits also extends to more challenging omni-modal tasks involving dual-modal inputs (e.g., vision and audio). Our experimental results with long reasoning traces further support the recent finding (rouditchenko2025omni) that it is possible to enhance audio content reasoning ability solely by injecting high quality reasoning text data. We provide a qualitative example in 8 on how reasoning traces on recognizing complex and compositional sound and speech events.
@@
A.5.2 Text-Only MMLU-PRO
MMLU-Pro (wang2024mmlupro) extends the original MMLU (hendrycks2021measuring) by incorporating more reasoning-intensive questions and enlarging the answer choice set. It covers 14 broad domains—including mathematics, physics, chemistry, and others—and consists of over 12,000 questions in total. In table 7 we show the extended table that was presented shorted in the main work.
A.5.3 Embodied Open-Ended QA Benchmark
We evaluate on a modified version of Needle in the Embodied Haystack (NiEH) (kim2025needlesembodiedhaystackenvironment), focusing on the single-evidence setting where one frame in a time-ordered trajectory suffices to answer the question. The test set contains 829 image–question pairs. In our evaluation, we restrict the input to a 2048-token window centered on the ground-truth (answer-bearing) frame, preserving temporal order so that multiple adjacent frames remain visible around the evidence.
@@ A.5.3 Embodied Open-Ended QA Benchmark
Table 8 summarizes results on the modified NiEH single-evidence benchmark (EM). Training on our data substantially improves performance: staged SFT+DPO achieves the best score (46.89), outperforming SFT only (39.67) and SFT+GRPO (35.06). All variants improve over the base model (30.67), underscoring the effectiveness of our data. Importantly, this task is entirely out-of-domain, and our training did not include open-ended questions.
@@
Table 10: Detailed cognitive behaviour statistics for various datasets. Values represent the average count of the behaviour on a subsample of 1000 examples of the dataset. Quantification methodology and terminology follows from  gandhi2025cognitivebehaviorsenableselfimproving; liao2025longperceptualthoughts.
Behaviour LongPerceptual Thoughts LongGrounded Thoughts (s1) LongGrounded Thoughts (s2) allenai/pixmo -ask-model DriveLMMo1 LENS SCI-Reason VLLA-Thinking Virgo Subgoal setting 0.036 0.12 0.55 0.017 0.235 0 0.032 0.17 0.24 Backtracking 0.354 0.35 0.68 0 0 0 0 0.37 0.79 Verification 0.26 0.33 0.76 0 0 0 0 0.45 0.89
@@
Figure 8: Temporal reasoning improvement illustrated by a qualitative example of a reasoning trace from the Qwen-2.5 Omni model post-trained on our data, compared to the base Qwen-2.5 Omni model, on an unseen audio reasoning question involving joint speaking and sound events.

### v2

@@ 5 Conclusion
We introduce a framework for synthesizing vision-centric reasoning data that focuses on scale, complexity, and cognitive richness, yielding over 1M high-quality examples. By leveraging grounded metadata and composition hardening, our pipeline generates diverse, verifiable visual problems and complex reasoning structures that significantly surpass prior efforts. Our experiments demonstrate that this data not only improves downstream performance on vision-centric benchmarks but also drives robust positive transfer across modalities and out-of-distribution tasks.
Impact Statements
Our work contributes to the democratization of visual reasoning capabilities through open datasets and reproducible training recipes. Since our data is synthesized from frontier models, potential risks include the propagation of existing biases found in the teacher models. We do not foresee immediate negative societal consequences beyond the general dual-use concerns applicable to any advancement in visual understanding/reasoning systems.
@@ A.1 Stage 1: Object-centric Problem Generation
Object-oriented metadata generation. Our data generation pipeline is composed of two primary stages: a scalable data generation stage focused on quantity and a compositional hardening stage focused on complexity. The first stage is illustrated in Figure 6. Our pipeline begins by leveraging a dense captioning model to produce a detailed textual description of the image. Concurrently, we process the image with Grounded SAM-2 (Ren et al., 2024) to extract a rich set of open-vocabulary bounding boxes and corresponding object tags, which we term Object Metadata. This metadata, including labels and coordinates (e.g., bag ($0.29,0.11,0.39,0.32$)), is integrated with the dense caption into a structured system prompt. This prompt guides a large language model (LLM Gen) to create object-centric MCQs that are inherently focused on specific visual elements. On average, each DOCCI image contains a median of 10.7 bounding tags, determined after applying a confidence cutoff of 0.9 from the detection source variable of the Florence-2 feature extractor (Xiao et al., 2024), a 0.7B vision encoder–decoder model, within the Grounded SAM-2 bounding box generation pipeline. To stabilize the bounding box–grounded MCQ generation process in the Stage 1 pipeline over diverse object-context, we constrain the maximum number of same-category instances per image to 9 (e.g., up to nine “tree” bounding box tags in a visual scene image).
Semantic filtering for question diversity. To ensure the quality and diversity of the generated data, we employ a two-step filtering process. First, an Qwen3-30B-A3B-Instruct-2507 (Yang et al., 2025) based LLM verifier assesses the factual correctness and logical soundness of each generated MCQ. Second, a semantic similarity filtering step is applied to discard redundant or overly similar questions, thereby enhancing the diversity of the final dataset.
For the semantic filtering, we represent each MCQ $i$ by a question stem embedding $q_{i}$, a selected answer embedding $a_{i}$, and an optional set of category tags $c_{i}$. Embeddings are derived from all-MiniLM-L6-v2, a 22M text embedding model in SentenceTransformer (Reimers and Gurevych, 2019), after standard text normalization. We define a composite similarity score between two MCQs, $i$ and $j$, as a weighted combination:
@@
A.2 Additional Analysis of Data Diversity
To investigate the source of the scalability gains reported in Section A.1, we analyze the semantic diversity of the generated supervision. We hypothesize that conditioning on dense captions alone leads to MCQ synthesis collapse, where the generator repeatedly targets the most salient image features (e.g., dominant objects, colors), creating redundant training signals. In contrast, we posit that conditioning on open-vocabulary bounding boxes (LGT) forces the generator to explore the long tail of visual concepts.
We test this by sampling $N=1000$ questions generated from three shared seed images for both LGT (Ours) and LPT (Baseline). We encode these questions using SentenceBERT (Reimers and Gurevych, 2019) all-MiniLM-L6-v2 and estimate their density in the semantic space using Gaussian Kernel Density Estimation (KDE).
Figure 7: Semantic Manifold Comparison. KDE contours (95% confidence) reveal that caption-only generation (LPT, right) suffers from mode collapse, clustering tightly around coarse visual features. Object-grounded generation (LGT, left) prevents this collapse, expanding the supervision manifold to cover finer-grained spatial and attribute reasoning.
As shown in Figure 7, the baseline distribution exhibits severe mode collapse, characterized by high density in a narrow semantic region. Quantitatively, LGT demonstrates superior coverage:
[…]
  • Information Density: LGT significantly reduces redundancy, with an average pairwise cosine similarity of 0.61 compared to 0.82 for LPT.
These results confirm that object-level metadata prevents MCQ synthesis collapse and maintains diversity at scale. By preventing saturation in the supervision signal, LGT enables the continuous performance scaling observed in our main results, whereas caption-based methods rapidly hit diminishing returns.
A.3 Vision-Centric Prompt Templates
When the previous LLM based text-to-text MCQs generation frameworks (Liao et al., 2025a) are highly relied on the image caption quality of the text, injection visual-grounded information has not been well explored. We carefully design a coordinate information guided prompt for the vision-centric MCQ generation with one main task instruction prompt, one core principle, and four auxiliary guidelines as shown in  A.3. We will open source the full data generation pipeline with the exact prompt instruction.
@@ A.3 Vision-Centric Prompt Templates
Prompt Task At the start of the process, we add a “computer vision expert” role-playing instruction; therefore, the text-LLM can better interpret bounding box coordinates and image tags while generating questions. In our experiments, using this “computer vision expert” instruction led to a noticeable improvement: the instruction-following F1 score rose from $96.3$ to $99.2$ during Stage 1 data generation. Similar role-playing (Zhang et al., 2025b), self-reflection (Shinn et al., 2023), and task-activating prompting (Yang et al., 2023) approaches have also been explored in recent work on domain-specific MCQ generation, such as in medical and speech understanding.
@@
Core Principle. To strengthen the model’s ability to reason over complex spatial and visual information in MCQs, we perform an error analysis on MTVQA (Tang et al., 2024) and MegaBench (Chen et al., 2024a) benchmarks that do not overlap with our test set, using the QwenVL-2.5 report as reference. The analysis reveals that most failure cases stem from two sources: 55.1% misunderstanding spatial relationships (e.g., identifying where an object mentioned in the question is located in the image) and 32.2% misinterpreting contextual information (e.g., inferring the function or color of the subject). Motivated by these observations, we establish four vision-centric principles to guide question construction, leveraging dense captions, bounding box labels, and coordinate information.
@@
In summary, LGT employs a carefully designed prompt instruction comprising $1,987$ tokens, which is substantially longer than the simpler vision-instruction prompt of $419$ tokens used in  (Liao et al., 2025a). The full prompt is provided in detail below.
@@
 Critical Reminders You must generate exactly {{ num_questions | default(4) }} questions. Questions MUST NOT disclose bounding box coordinates or specific object labels. Use generic terms like "the object," "the item," or "the element" instead of "{{ bbox_label }}". Answers MUST INCLUDE the exact object label and integer coordinates in the specified format: "{{ bbox_label }}, [{{ bbox_coordinates[0]|round|int }}, {{ bbox_coordinates[1]|round|int }}, ...], [specific answer]". Questions must focus on visual properties or relationships of the object within the specified bounded region, requiring careful inspection.
A.4 Stage 2: Compositional Problem Generation
@@
Simple CoT VLM distillation and Thought Expansion A core challenge in synthesizing reasoning traces is that open-weight VLMs often lack the rich, non-linear cognitive behaviors (e.g., subgoal setting, backtracking) seen in frontier LLMs, while human annotation is prohibitively expensive at scale. To address this, we adopt the two-step distillation process from LPT, as illustrated in Figure 1 (bottom). The prompt templates used for this distillation process are adapted from a baseline in (Liao et al., 2025a).
@@
A.5 Local Qwen CoT Verification Prompt
@@
A.6 Additional Experimental Results
A.6.1 Multimodal Audio Understanding and Omni Model
We emphasize our data is completely vision-centric. Here, we conduct experiments on out-of-domain (OOD) complex audio reasoning tasks using the Qwen2.5-Omni-7B-Instruct (Xu et al., 2025) model on the Multimodal Audio Understanding (MMAU) benchmark (Sakshi et al., 2024). Our methodology leverages the modular architecture of Qwen-Omni, which separates the reasoning component (“thinker”) from the answer generation component (“talker”). We first fine-tune only the “thinker” dense module using our synthesized data. We keep our modality-specific audio encoder frozen when tuning the backbone dense model. After this stage, the fine-tuned “thinker” is merged back with the original, pre-trained “talker” module. This strategy aims to enhance the model’s intrinsic reasoning capabilities without directly altering the audio-specific knowledge contained within the “talker” module, thereby mitigating catastrophic forgetting.
@@ A.6.1 Multimodal Audio Understanding and Omni Model
Training on 500k LGTs already pushes the model’s performance to 71.80, outperforming the other leading omni-models. By scaling to 1M LGTs, our model achieves a new state-of-the-art average score of 72.32 among the tested configurations. These gains are driven by significant improvements in non-speech acoustic reasoning, with the MMAU-Sound score rising from 76.77 to 78.30 and the MMAU-Music score increasing from 67.33 to 70.20. This result underscores that enhancing core reasoning abilities with our vision-centric data can positively transfer to and improve performance on OOD audio tasks. This complex tracing of learning benefits also extends to more challenging omni-modal tasks involving dual-modal inputs (e.g., vision and audio). Our experimental results with long reasoning traces further support the recent finding (Rouditchenko et al., 2025) that it is possible to enhance audio content reasoning ability solely by injecting high quality reasoning text data. We provide a qualitative example in 9 on how reasoning traces on recognizing complex and compositional sound and speech events.
@@
A.6.2 Text-Only MMLU-PRO
MMLU-Pro (Wang et al., 2024) extends the original MMLU (Hendrycks et al., 2021) by incorporating more reasoning-intensive questions and enlarging the answer choice set. It covers 14 broad domains (i.e., mathematics, physics, chemistry, and others) and consists of over 12,000 questions in total. In table 7 we show the extended table that was presented shorted in the main work.
A.6.3 Embodied Open-Ended QA Benchmark
We evaluate on a modified version of Needle in the Embodied Haystack (NiEH) (Kim and Ammanabrolu, 2025), focusing on the single-evidence setting where one frame in a time-ordered trajectory suffices to answer the question. The test set contains 829 image–question pairs. In our evaluation, we restrict the input to a 2048-token window centered on the ground-truth (answer-bearing) frame, preserving temporal order so that multiple adjacent frames remain visible around the evidence.
@@ A.6.3 Embodied Open-Ended QA Benchmark
Table 8 summarizes results on the modified NiEH single-evidence benchmark (EM).
@@
A.7 Experiments on MathVision and VisuLogic Reasoning
In Figure 9, we additionally evaluated our best model finetuned on our data on MathVision and VisuLogic. Remarkably, despite our data containing no math questions, we see meaningful improvements on both benchmarks. This validates our data’s generalization capability and demonstrates that strengthening core reasoning abilities transfers to out-of-domain tasks.
Table 9: Our best model finetuned on our data on MathVision and VisuLogic. Despite our data containing no math questions, we see meaningful improvements on both benchmarks.
Model MathVision VisuLogic
Qwen2.5-VL-7B-Instruct 49.57 11.5
Ours (SFT and RL) 51.77 (+2.2) 19.3 (+7.8)
@@ A.7 Experiments on MathVision and VisuLogic Reasoning
LPT BaseModel 750K None None 0.682
Ours BaseModel 750K None None 0.716
LPT BaseModel 0 GRPO 10K 0.695
50K 0.683
100K 0.669
LPT SFT 750K GRPO 10K 0.662
50K 0.685
100K 0.709
Ours BaseModel 0 GRPO 70K 0.704
SFT 750K DPO 70K 0.737
SFT 750K DPO 129K 0.740
SFT 750K GRPO 70K 0.757
SFT 750K GRPO 129K 0.746
Table 12: Impact of CoT Richness (Stage 1 Data). Simple traces (Short CoT) hurt performance; only rich, expanded traces (Full CoT) yield best results.
Configuration CV-Bench V*Bench
Base (Qwen2.5-VL-7B) 0.740 0.485
+ Thinking Prompt (No SFT) 0.754 0.551
+ SFT (No CoT) 0.787 0.583
+ SFT (Short CoT) 0.714 0.540
+ SFT (Full CoT - Ours) 0.813 0.597
Table 13: Detailed cognitive behaviour statistics for various datasets. Values represent the average count of the behaviour on a subsample of 1000 examples of the dataset. Quantification methodology and terminology follows from  (Gandhi et al., 2025; Liao et al., 2025a).
Behaviour LongPerceptual Thoughts LongGrounded Thoughts (s1) LongGrounded Thoughts (s2) allenai/pixmo -ask-model DriveLMMo1 LENS SCI-Reason VLLA-Thinking Virgo Subgoal setting 0.036 0.12 0.55 0.017 0.235 0 0.032 0.17 0.24 Backtracking 0.354 0.35 0.68 0 0 0 0 0.37 0.79 Verification 0.26 0.33 0.76 0 0 0 0 0.45 0.89 Sum 0.65 0.8 1.99 0.017 0.235 0 0.032 0.99 1.92
@@
Figure 9: Temporal reasoning improvement illustrated by a qualitative example of a reasoning trace from the Qwen-2.5 Omni model post-trained on our LGT data, compared to the base Qwen-2.5 Omni model, on an unseen audio reasoning question involving joint speaking and sound events recognition.